A student wants to estimate the mean score of all college students for a particular exam. First use the range rule of thumb to make a rough estimate of the standard deviation of those scores. Possible scores range from 600 to 1400 Use the TI-83/84 and the estimated standard deviation to determine the sample size corresponding to a 95% confidence level and a margin of error of 100 points.

A confidence level of 95% requires a minimum sample size of____.
Round UP to the nearest integer. 
Answer #1

I have answered the question below

Please up vote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

zcrit = 1.96

Standard deviation = Range/4

Population SD = sigma = 1400-600/4 = 200

E = 100

=> size = (1.96*200/100)2 = 15.36 ~ 16 samples

Note: Please try with 15 samples if the above solution comes as incorrect